For each of the investments shown in the following table, calculate the rate of return earned...

  1. For each of the investments shown in the following table, calculate the rate of return earned over the specified time period.

Investment       Cash Flow during the Period      Beginning Value                        Ending Value

     A                  $15,000                                    $120,000                      $118,000

     B                  $80                                          $600                             $500

     C                  $7,000                                      $45,000                        $48,000

  1. Assume that the risk-free rate is currently 9% and the market return is currently 13%.

Calculate the required return for an asset with the following beta:

Investment       Beta                 Required Return

Asset A             .8                     

Asset B             1.0

Asset C             1.5

Asset D             2.0
Answer #1

Return=(Ending Value+Cash flow)/Beginning Value-1

1.
=(118000+15000)/120000-1
=10.83%

2.
=(500+80)/600-1
=-3.33%

3.
=(48000+7000)/45000-1
=22.22%

Required return=risk free rate+beta*(market return-risk free rate)

1.
=9%+0.8*(13%-9%)
=12.20%

2.
=9%+1*(13%-9%)
=13.00%

3.
=9%+1.5*(13%-9%)
=15.00%

4.
=9%+2*(13%-9%)
=17.00%